OpenStudy (anonymous):

Solve 10cosx + 5 = 15

OpenStudy (anonymous):

get the similar items together

OpenStudy (anonymous):

10cosx = 15-5

OpenStudy (anonymous):

and then it is simple from here

OpenStudy (anonymous):

10cosx+5=15 10cosx=15-5 10cosx=10 divide both sides by 10 cosx=1 x=arc cos 1 x=0